write the slope-intercept equation for the line that passes through (-8,11) and is perpendicular to -2x-7y=14
Answer by mananth(16949)

-2x-7y =14
Find the slope of this line
-7y=2x-14
Divide by -7
y =- 2/7x+2
Compare this equation with y=mx+b
slope m =-0.29

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be 3.50
3.5=
m=3.50 ,point (-8,11)
Find b by plugging the values of m & the point in
y=mx+b
11=-28+b
b=39
m=3.5
The required equation is y=3.5x +39